# Tuning the magnetic properties of spin-split antiferromagnet MnTe through pressure

### v1

@@ III Discussion
The dependence of the magnetic moment and $T_{\mathrm{N}}$ on pressure is likely to have interesting applications. For example, chemical pressure induced by replacing Te with the smaller chalcogenides Se or S may mimic the effect of physical pressure and raise $T_{\mathrm{N}}$ higher above room temperature, thereby expanding the temperature range available for practical applications while retaining other properties of MnTe. Indeed, an earlier report Chehab et al. (1986) indicates that $T_{\mathrm{N}}$ rises as expected when up to 20% of the Te is replaced with Se. It may also be possible to induce a negative chemical pressure to lower $T_{\mathrm{N}}$ or increase the low-temperature ordered moment. The distinct VB- and CB-spin-splitting behavior with pressure likewise invites future experimental study and may enable novel applications in spintronics. Additional investigations will be necessary to explore these possibilities fully. Overall, this study provides a promising outlook on the use of pressure to optimize the properties of MnTe and potentially other NRSS antiferromganets.
IV Methods
IV.1 Neutron diffraction procedure and analysis
The polycrystalline MnTe sample was made by reacting the elements in an alumina crucible sealed inside an evacuated silica ampoule as described in Ref. Moseley et al. (2022). The sample was stable in air; no additional impurity was observed after one month of air exposure.
Neutron powder diffraction was performed at Oak Ridge National Laboratory’s HB-2A POWDER beamline with an incident wavelength of 2.41 Å. The sample was loaded into a Teflon container which in turn was loaded into a piston cylinder cell consisting of a main Al cylinder with a prestressed CuBe insert as well as a WC piston assembly with optical access for fluoresence measurements Podlesnyak et al. (2018). Fluorinert was used as pressure medium as a prior test revealed minimal evaporation under the temperature conditions used. Further, fluorinert remains hydrostatic under the pressure conditions used here Varga et al. (2003). This cell was chosen to minimize background contamination at MnTe’s magnetic Bragg peaks. Neutron diffraction patterns were collected at applied pressures of 0, 0.3, 0.5, and 1 GPa, where we used ruby fluorescence and its associated pressure equation of state Dewaele et al. (2008) to measure these pressures. Using the equation of state found in Ref. Mimasaka et al. (1987) and lattice parameters from our Rietveld refinements, we find that our pressures are accurate within 0.1 GPa. For each pressure, we used the cryostat at HB2A to acquire the full powder diffraction pattern at 250 and 350 K. We also acquired full powder diffraction patterns for at least one pressure at $T=$280, 295, 320, and 370 K. In these diffraction patterns the reflections 001, 100, 002 and 101 are well resolved and separated from any pressure cell scattering. However, most other peaks were heavily contaminated by the CuBe pressure cell background, so we had a limited $2\theta$ range $(17^{\circ}<2\theta<~{}55^{\circ})$ to do Rietveld refinements.
Rietveld refinements were performed using FullProf Arcelus et al. (2024). We started by fitting lattice parameters and a preferred orientation parameter to each of the patterns collected at 350 K. Because $T_{N}<350$ K for all pressures, the observed peaks from MnTe are purely nuclear with no magnetic contribution. The preferred orientation parameter $G_{1}$ was nearly identical for fits. Then, for all other temperatures, we fixed the preferred orientation parameter for each pressure to the values found at 350 K. $B_{\mathrm{iso}}$ was fixed to 1.0 Å2 for all fits, since the limited usable $2\theta$ range imposed by the pressure cell precluded reliable refinements of this parameter. The lattice parameters and ordered moment were then refined for all lower-temperature data sets.
We also conducted an order parameter (OP) scan at each pressure. The OP scans measured the intensity of MnTe’s antiferromagnetic 001 peak. This allows us to precisely determine the Néel temperature ($T_{\mathrm{N}}$) at each applied pressure, because above $T_{\mathrm{N}}$, the magnetic contribution to the Bragg peaks vanishes. For 0 GPa, we scanned from 250 K to 350 K in steps of 2 K, for 1 GPa we scanned from 270 K to 370 K in steps of 1 K, and for the other pressures we scanned from 270 K to 350 K in steps of 1 K.
IV.2 DFT calculations of electronic properties
We used a plane-wave pseudopotential density functional theory (DFT) method to perform electronic calculations as implemented by the Vienna Ab initio Simulation Package (VASP) Kresse and Furthmüller (1996a, b). We utilized the projector augmented-wave (PAW) method Blöchl (1994); Kresse and Joubert (1999) with the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional within the generalized gradient approximation (GGA) Perdew et al. (1996) plus an on-site potential term $U$ of 4 eV taken from Ref. Alaei et al. (2025) to correct the spurious self-interaction errors in Mn-$d$ orbitals. The PAW-PBE pseudopotentials are adopted with the valence electrons in MnTe including 13 electrons in Mn (2p64s23d5) and 6 electrons in Te (5s25p4). We set the following parameters in our DFT calculations: (i) The cut-off total energy for the plane-wave basis was 360 eV; (ii) The convergence tolerance for total energy was $10^{-6}$ eV; (iii) The convergence tolerance of relaxation was $10^{-2}$ eV/Å; (iv) The $\Gamma$-centered $k$-point sampling grid was $5\times 5\times 3$ for relaxation and static self-consistent calculation.
Data Availability The datasets generated and/or analysed during the current study are available in Zenodo with the identifier https://doi.org/10.5281/zenodo.15353048.
@@ IV.2 DFT calculations of electronic properties
Acknowledgements We thank Malcolm Cochran of Oak Ridge National Laboratory for valuable assistance selecting the best pressure medium for the experiment, as well as Xiuwen Zhang for valuable discussions. The neutron scattering work carried out by EPC and BAF was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences (DOE-BES) through Award No. DE-SC0021134. The DFT calculations of electronic properties performed by AZ and J-XX were supported by DOE-BES Materials Sciences and Engineering Division under Award No. DE-SC0010467. The DFT calculations used resources of the National Energy Research Scientific Computing Center, which is supported by the Office of Science of the U.S. Department of Energy. J-XX used the VASP software under the license from Alex Zunger. A portion of this research used resources at the High Flux Isotope Reactor, a DOE Office of Science User Facility operated by the Oak Ridge National Laboratory. The beam time was allocated to HB2A on proposal number IPTS-32506. Sample synthesis and characterization at ORNL was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division.

### v2

@@ IV Discussion
The dependence of the magnetic moment and $T_{\mathrm{N}}$ on pressure is likely to have interesting applications. For example, chemical pressure induced by replacing Te with the smaller chalcogenides Se or S may mimic the effect of physical pressure and raise $T_{\mathrm{N}}$ higher above room temperature, thereby expanding the temperature range available for practical applications while retaining other properties of MnTe. Indeed, an earlier report [54] indicates that $T_{\mathrm{N}}$ rises as expected when up to 20% of the Te is replaced with Se. It may also be possible to induce a negative chemical pressure to lower $T_{\mathrm{N}}$ or increase the low-temperature ordered moment. The distinct VB- and CB-spin-splitting behavior with pressure likewise invites future experimental study and may enable novel applications in spintronics. Additional investigations will be necessary to explore these possibilities fully. Overall, this study provides a promising outlook on the use of pressure to optimize the properties of MnTe and potentially other NRSS antiferromganets.
Data Availability The data that support the findings of this article are openly available [55].
@@
Acknowledgements We thank Malcolm Cochran of Oak Ridge National Laboratory for valuable assistance selecting the best pressure medium for the experiment, as well as Xiuwen Zhang for valuable discussions. The neutron scattering work carried out by EPC and BAF was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences (DOE-BES) through Award No. DE-SC0021134. The DFT calculations of electronic properties performed by AZ and J-XX were supported by DOE-BES Materials Sciences and Engineering Division under Award No. DE-SC0010467. The DFT calculations used resources of the National Energy Research Scientific Computing Center, which is supported by the Office of Science of the U.S. Department of Energy. J-XX used the VASP software under the license from Alex Zunger. A portion of this research used resources at the High Flux Isotope Reactor, a DOE Office of Science User Facility operated by the Oak Ridge National Laboratory. The beam time was allocated to HB-2A on proposal number IPTS-32506. Sample synthesis and characterization at ORNL was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division.